OpenStudy (anonymous):

simplify the rational expression 2x^2 -4x -6 / x^2 +x -12

jimthompson5910 (jim_thompson5910):

(2x^2 -4x -6)/(x^2 +x -12) (2(x+1)(x-3))/(x^2 +x -12) (2(x+1)(x-3))/((x+4)(x-3)) (2(x+1))/(x+4) (2x+2)/(x+4) So \[\large \frac{2x^2 -4x -6}{x^2 +x -12}=\frac{2x+2}{x+4}\]